# SAKE: Towards Editing Auditory Attribute Knowledge of Large Audio-Language Models

### v1

@@ 6 Conclusions
In this paper, we present the first study on editing auditory attribute knowledge in LALMs. We introduce SAKE, a benchmark designed to evaluate this type of knowledge editing across four dimensions. Through comprehensive experiments applying seven representative methods to two state-of-the-art LALMs, we reveal limitations in preserving non-target auditory knowledge and in generalizing edits to interconnected world knowledge during reasoning. Our findings provide new insights into this unexplored direction and establish a foundation for future research on developing more robust and specialized knowledge editing methods for auditory modalities.
Limitations. As the first benchmark for editing auditory attributes in LALMs, we focus on four representative attributes, though expanding further would allow broader evaluation. Similarly, while using two models follows common practice (Cheng et al., 2023; Zhang et al., 2024), including more LALMs could yield deeper insights. In addition, while our study focuses on speech-to-text LALMs, future work could extend to speech-to-speech LALMs (Fang et al., 2025; Xie & Wu, 2024; Yang et al., 2024b; Lin et al., 2024a; Chiang et al., 2025), which are likely to present greater challenges. We leave these directions for future work.
Ethics Statement
This work studies knowledge editing in large audio-language models (LALMs). Our datasets are publicly available and licensed for research. While knowledge editing offers benefits such as correcting errors and mitigating bias, we acknowledge potential misuse (e.g., manipulation of model knowledge) and emphasize that our contributions are intended for socially responsible applications. We report our methods and results transparently to ensure reproducibility.
Reproducibility Statement
Our implementation of editing methods is built on the EasyEdit toolkit (Wang et al., 2024b), with hyperparameters provided in Appendix C.2. To reduce randomness in the LLM-as-a-judge evaluation, we fix the temperature at 0. We also report the version of the judge model and include the evaluation prompts in Appendix D. Together, these details help ensure the reproducibility.
@@ Appendix A The Usage of Large Language Models (LLMs)
Appendix B Dataset Statistics
We report the dataset statistics in Table 2 and Table 3, respectively, for the training and testing datasets, highlighting the diversity of the data.
@@ Appendix B Dataset Statistics
((a)) Question Length (word).
@@
((b)) Audio/speech duration (s).
@@
((a)) Question Length (word).
@@
((b)) Audio/speech duration (s).
@@
Appendix C Implementation Details of the Editing Methods
C.1 Editing Methods
Fine-tuning adapts the model via gradient descent on selected components. We apply it to the last layer of the LLM backbone and the modality connector between audio encoders and the backbone in LALMs.
Knowledge Editor (KE) (De Cao et al., 2021) employs a hyper-network to update parameters. It leverages a bidirectional LSTM with constrained optimization to predict weight updates.
MEND (Mitchell et al., 2022) uses a hypernetwork to generate parameter updates by decomposing fine-tuning gradients into low-rank forms and transforming them into parameter updates.
C.2 Implementation Details
The hyperparameters of each editing method are shown in Table 4, and their approximate execution time on NVIDIA H100 GPU is reported in Table 5. For FT (LLM) and FT (Audio), early stopping is applied when the loss falls below 1e-2. For KE and MEND, which require training a hyper-network, we train the hyper-network for one epoch due to its rapid convergence on the validation set. For UnKE, the editing process consists of two stages with separate hyperparameters. The first stage uses v_step iterations to find the desired key vector, with its own learning rate v_LR. The second stage uses optim_step iterations to update the weights of the chosen layer so that it naturally produces the new key vector, controlled by a separate learning rate optim_LR. To prevent interference with unrelated knowledge, stage 2 also incorporates preservation data. Specifically, we use 20 preservation samples for text and 5 for audio. For I-IKE and IE-IKE, the edited knowledge is encoded through natural language instructions in the system prompt, using the prompt template provided below. In IE-IKE, the demonstration examples are retrieved from the training set according to the editing pair. For example, when editing the model’s perception of speaker emotion from happy to sad, we retrieve an editing instance from the training set involving the same (happy, sad) editing pair. The demonstrating examples cover reliability, generality, and locality.
System Prompt Template for I-IKE and IE-IKE You are performing in-context knowledge editing, but only for questions about audio inputs. From now on, when analyzing any audio clip, you must apply the following edits consistently: EDITS (New facts): {edits} where each entry has the form pre_edit $\rightarrow$ post_edit. Rules: 1. If your reasoning or prediction about an audio clip would normally lead to pre_edit, you must instead treat it as post_edit. 2. All properties, attributes, and facts that belong to post_edit must be applied consistently, as if the audio were actually from post_edit. 3. If the user’s question is unrelated to these edits, you should answer normally without making changes. 4. Always ensure your final answers are fully consistent with the edited mapping.
@@ C.2 Implementation Details
DeSTA2.5-Audio 15 31st layer of Transformer Module Adam 1e-5
Qwen2-Audio 15 31st layer of Transformer Module Adam 1e-4
@@
DeSTA2.5-Audio 15 perception.connector Adam 1e-4
Qwen2-Audio 15 multi_modal_projector Adam 1e-4
@@
DeSTA2.5-Audio 1 layer 29. 30, 31 of Transformer Module RMSprop 3e-4
Qwen2-Audio 1 layer 29. 30, 31 of Transformer Module RMSprop 3e-4
@@
DeSTA2.5-Audio 1 layer 29. 30, 31 of Transformer Module Adam 1e-6
Qwen2-Audio 1 layer 29. 30, 31 of Transformer Module Adam 1e-6
@@
DeSTA2.5-Audio 25/50 layer 15 of Transformer Module 20(text)/5(audio) 5e-1/2e-4
Qwen2-Audio 25/50 layer 20 of Transformer Module 20(text)/5(audio) 5e-1/2e-4
@@
FT (LLM) DeSTA2.5-Audio - 5h 50m 2h 0m
Qwen2-Audio - 2h 35m 1h 15m
FT (Audio) DeSTA2.5-Audio - 4h 20m 2h 0m
Qwen2-Audio - 3h 0m 1h 30m
KE DeSTA2.5-Audio 4h 10m 8h 40m 9h 45m
Qwen2-Audio 2h 10m 4h 50m 3h 20m
MEND DeSTA2.5-Audio 4h 10m 4h 0m 4h 10m
Qwen2-Audio 2h 20m 2h 10m 1h 50m
UnKE DeSTA2.5-Audio - 3h 0m 1h 05m
Qwen2-Audio - 1h 35m 0h 30m
I-IKE DeSTA2.5-Audio - 6h 0m 3h 45m
Qwen2-Audio - 3h 0m 2h 0m
IE-IKE DeSTA2.5-Audio - 3h 20m 8h 0m
Qwen2-Audio - 2h 0m 3h 50m
Appendix D More Details and Evaluation Prompts for LLM-as-a-judge
@@
We use GPT-5 mini with minimal reasoning effort as the judge in our LLM evaluator, balancing strong alignment with human judgments and cost efficiency. The temperature is set to 0 to ensure maximal reproducibility. Human verification was conducted on 210 randomly selected samples for correctness and consistency respectively, yielding 98.57% and 97.62% agreement. As described in Sec. 4.3, evaluation tasks are categorized into correctness and consistency, with prompts provided below. To further reduce costs, some cases are judged directly without invoking the LLM evaluator: outputs that are empty are marked as incorrect/inconsistent, and responses that exactly match the ground truth (in correctness) or the original model’s output (in consistency) are marked as correct/consistent.
@@
Ground truth answer: {ground_truth} Model-generated response: {model_response} Return format.
@@
Original Output: {original_output} New Output: {new_output} Task. Decide if these two outputs are semantically consistent. Outputs are consistent if they express the same main message, even with different wording, phrasing, or level of detail. Outputs are inconsistent if they contradict each other, reach different conclusions, or provide fundamentally different answers. Return format.
@@
Table 6: Detailed results of the four metrics of each auditory attribute across different editing methods on DeSTA2.5-Audio under single editing. Attr. denotes auditory attributes, and Port. denotes portability. For generality and audio locality, Avg. indicates the average performance across all types of the corresponding metric. (%)
@@
FT (LLM) ALL 99.75 98.75 99.67 99.00 97.58 65.11 88.08 15.56 74.67 69.75 82.58 19.42
Animal 99.33 99.11 99.67 99.33 98.33 55.08 90.67 9.00 56.33 64.33 82.00 20.00
Emotion 100.00 99.56 99.33 100.00 99.33 53.25 81.67 7.00 59.67 64.67 79.33 23.00
Gender 99.67 96.78 99.67 96.67 94.00 84.67 90.33 - 88.33 75.33 86.33 7.00
Language 100.00 99.56 100.00 100.00 98.67 72.33 89.67 30.67 94.33 74.67 82.67 27.67
FT (Audio) ALL 99.50 86.14 96.75 84.92 76.75 68.09 78.67 48.11 76.17 64.42 100.00 55.50
Animal 99.00 85.78 97.33 83.00 77.00 62.00 81.67 40.67 60.00 65.67 100.00 30.33
Emotion 99.67 78.67 95.33 73.67 67.00 60.92 77.67 38.67 66.67 60.67 100.00 46.33
Gender 99.67 92.33 99.33 92.67 85.00 77.67 78.33 - 87.33 67.33 100.00 82.33
Language 99.67 87.78 95.00 90.33 78.00 74.17 77.00 65.00 90.67 64.00 100.00 63.00
KE ALL 99.58 99.17 99.25 99.25 99.00 79.29 96.33 43.89 76.33 91.75 92.50 18.42
Animal 98.67 97.33 97.67 97.33 97.00 77.33 97.00 61.67 59.00 91.67 91.33 23.67
Emotion 100.00 100.00 100.00 100.00 100.00 66.17 98.33 16.33 59.33 90.67 93.33 16.00
Gender 99.67 99.56 99.33 99.67 99.67 93.89 96.33 - 92.33 93.00 92.67 9.33
Language 100.00 99.78 100.00 100.00 99.33 83.42 93.67 53.67 94.67 91.67 92.67 24.67
MEND ALL 95.33 95.00 95.83 95.17 94.00 71.47 93.42 17.22 74.67 87.00 92.08 19.42
Animal 90.00 89.44 91.33 89.00 88.00 63.50 93.67 19.33 57.00 84.00 92.00 20.33
Emotion 99.67 98.44 99.00 99.00 97.33 62.25 94.33 9.00 59.67 86.00 94.00 27.33
Gender 98.67 97.00 97.33 99.33 94.33 91.89 97.33 - 88.33 90.00 90.00 4.33
Language 93.00 95.11 95.67 93.33 96.33 73.33 88.33 23.33 93.67 88.00 92.33 25.67
UnKE ALL 96.30 89.28 89.67 94.25 83.92 56.04 67.67 10.56 71.92 62.67 87.38 18.73
Animal 99.30 95.67 98.00 98.33 90.67 51.83 69.33 15.67 59.67 62.67 86.00 27.30
Emotion 96.00 91.56 91.33 94.33 89.00 50.42 66.33 12.00 57.00 66.33 87.60 20.00
Gender 99.60 86.11 85.67 99.67 73.00 72.33 58.67 - 94.33 64.00 86.60 7.30
Language 90.30 83.78 83.67 84.67 83.00 53.67 76.33 4.00 76.67 57.67 89.30 20.30
I-IKE ALL 73.00 61.47 64.67 59.58 60.17 65.40 79.42 50.67 73.67 54.17 62.08 71.42
Animal 90.33 92.89 92.67 93.00 93.00 54.75 84.33 22.00 59.67 53.00 60.00 78.33
Emotion 59.33 50.89 58.00 47.00 47.67 60.08 80.67 45.33 62.33 52.00 59.00 65.33
Gender 69.33 64.33 71.33 60.33 61.33 73.33 81.67 - 79.67 58.67 64.33 73.00
Language 38.67 37.78 36.67 38.00 38.67 75.42 71.00 84.67 93.00 53.00 65.00 69.00
IE-IKE ALL 40.58 39.61 41.50 38.67 38.67 58.76 70.08 49.89 70.75 42.08 56.25 34.58
Animal 77.00 78.56 80.00 76.67 79.00 51.17 78.67 23.33 59.00 43.67 60.33 44.33
Emotion 27.00 26.67 28.67 25.33 26.00 51.33 68.00 33.33 61.33 42.67 51.33 29.00
Gender 41.67 37.67 43.00 36.00 34.00 61.67 74.67 - 67.67 42.67 60.33 32.67
Language 16.67 15.56 14.33 16.67 15.67 71.58 59.00 93.00 95.00 39.33 53.00 32.33
@@
Table 7: Detailed results of the four metrics of each auditory attribute across different editing methods on Qwen2-Audio under single editing. Attr. denotes auditory attributes, and Port. denotes portability. For generality and audio locality, Avg. indicates the average performance across all types of the corresponding metric. (%)
Method Attr. Relability Generality Audio Locality Text Locality Port.

### v2

@@ 6 Conclusions
We present the first study on editing auditory attribute knowledge in LALMs and introduce SAKE, the first benchmark for evaluating this capability across four dimensions. Experiments with eight methods on three LALMs reveal challenges in preserving non-target auditory knowledge and propagating edits during reasoning. Under sequential editing, most methods suffer severe forgetting or degeneration. Overall, fine-tuning the modality connector serves as a strong and more balanced baseline for future work. In contrast, many methods that are effective in textual domains degrade substantially on auditory attributes, highlighting the unique challenges of auditory knowledge editing.
Impact Statement
This work establishes auditory attribute knowledge editing as a previously unexplored research direction for large audio-language models (LALMs). By systematically examining the modification of perceptual auditory knowledge, we identify fundamental limitations of existing editing methods. We introduce SAKE, the first benchmark for evaluating this capability, and present a comprehensive analysis that highlights key challenges in edit reliability, knowledge generalization, preservation, and propagation. We expect SAKE to serve as a foundation for future research in this area. All related resources will be publicly released after the review process.
Knowledge editing can enhance LALMs through error correction, bias mitigation, and personalization. At the same time, modifying perceptual knowledge introduces risks of misuse, including adversarial manipulation of model behavior. This work aims to promote the responsible development and deployment of LALMs by explicitly characterizing these limitations and informing future research on principled model editing.
@@ Appendix A The Usage of Large Language Models (LLMs)
Appendix B Dataset Details
B.1 Audio Sources
Audios are primarily sourced from the SAKURA benchmark (Yang et al., 2025b), which compiles data from CommonVoice (Ardila et al., 2020), CREMA-D (Cao and others, 2014), ESC-50 (Piczak, 2015), and Animal-Sound Dataset (Şaşmaz and Tek, 2018). In addition, we supplement SAKURA with a subset of samples drawn from Dynamic-SUPERB Phase-2 (Huang et al., 2025) for type 4 audio locality data.
B.2 Dataset Statistics
We summarize the more details about the statistics of the training and testing sets in Tables 4 and 5, highlighting the data diversity.
@@ B.2 Dataset Statistics
(a) Question Length (word).
@@
(b) Audio/speech duration (s).
@@
(a) Question Length (word).
@@
(b) Audio/speech duration (s).
@@
Appendix C Design Choices
C.1 Attributes
SAKE covers various auditory attributes, namely animal sound, speaker emotion, speaker gender, and spoken language, because they are fundamental to speech and audio understanding, widely used in real-world applications, and consistently evaluated in existing LALM benchmarks (Sakshi et al., 2025; Yang et al., 2025b; Huang et al., 2025; Yang et al., 2025c). The attribute label sets are provided as follows:
  • Gender: Male, Female
  • Emotion: Happy, Disgust, Sad, Fear, Angry
  • Language: English, German, Spanish, French, Italian, Chinese, Japanese, Korean
  • Animal sound: Dog, Cat, Pig, Cow, Frog, Hen, Rooster, Sheep, Crow
As the first study on auditory knowledge editing, focusing on these core attributes provides a clear and representative starting point. Moreover, these attributes are supported by comparatively rich resources, including curated audio datasets and question–answer pairs, which are essential for constructing reliable evaluations, particularly for our portability track that involves multi-hop reasoning (Yang et al., 2025b). This selection establishes a focused and principled foundation for the first benchmark in this area.
C.2 Models
We select Qwen2-Audio (Chu et al., 2024), DeSTA2.5-Audio (Lu et al., 2025b), and Audio Flamingo 3 (Ghosh et al., 2025a) as strong, widely recognized, and representative open-source LALMs. Qwen2-Audio is widely adopted as a community baseline in studies on reasoning (Sakshi et al., 2025; Yang et al., 2025b; Ma et al., 2025b), safety (Yang et al., 2025d; Feng et al., 2025), and interpretability (Yang et al., 2025a). DeSTA2.5-Audio and Audio Flamingo 3 are more recent models that achieve competitive performance across multiple benchmarks (Sakshi et al., 2025; Huang et al., 2025; Ma et al., 2025b; Wang et al., 2025), reflecting recent advances in LALMs.
Importantly, prior work (Yang et al., 2025b) shows that these models already possess the world knowledge required by our portability dataset, whose connected attribute knowledge is drawn from the SAKURA benchmark. Moreover, they demonstrate state-of-the-art multi-hop reasoning that integrates auditory understanding with world knowledge. This allows observed portability errors to be attributed to failures in edit propagation, rather than limitations in world knowledge or inherent reasoning capability.
As our goal is to evaluate the editing methods rather than to compare LALMs themselves, we follow standard practice in knowledge editing research by applying each method to two to three representative models (Cheng et al., 2023; Du. et al., 2025; Zhang et al., 2025; Ma et al., 2025a; Deng et al., 2025; Wang et al., 2024b; Meng et al., 2022; Chen et al., 2025). Importantly, the selected models belong to different families and differ substantially in their audio encoders, LLM backbones, modality adaptation mechanisms, and training strategies, thereby covering a diverse range of LALM designs. The fact that our key observations consistently emerge across these models and across a wide range of editing methods suggests that the identified challenges stem from limitations of current editing techniques, rather than from idiosyncrasies of any particular model. This supports the suitability of our model selection as a testbed for evaluating auditory knowledge editing.
Appendix D Implementation Details of the Editing Methods
Our implementation of editing methods is built on the EasyEdit toolkit (Wang et al., 2024c).
@@ Appendix D Implementation Details of the Editing Methods
DeSTA 15 Layer 31 of Transformer Module Adam 1e-5
Qwen 15 Layer 31 of Transformer Module Adam 1e-4
AF 15 Layer 27 of Transformer Module Adam 2e-5
@@
DeSTA 15 perception.connector Adam 1e-4
Qwen 15 multi_modal_projector Adam 1e-4
AF 15 multi_modal_projector Adam 5e-4
@@
DeSTA 1 layer 29, 30, 31 of Transformer Module RMSprop 3e-4
Qwen 1 layer 29, 30, 31 of Transformer Module RMSprop 3e-4
AF 1 layer 25, 26, 27 of Transformer Module RMSprop 3e-4
@@
DeSTA 1 layer 29, 30, 31 of Transformer Module Adam 1e-6
Qwen 1 layer 29, 30, 31 of Transformer Module Adam 1e-6
AF 1 layer 25, 26, 27 of Transformer Module Adam 1e-6
@@
DeSTA 25/50 layer 15 of Transformer Module 20(text)/5(audio) 5e-1/2e-4
Qwen 25/50 layer 20 of Transformer Module 20(text)/5(audio) 5e-1/2e-4
AF 25/50 layer 20 of Transformer Module 20(text)/5(audio) 5e-1/2e-4
WISE
Model $\alpha$, $\beta$, $\gamma$, act_ratio Edit Layer preserve_data edit_LR/n_iter
DeSTA 2,20,10,0.88 layer 29 of Transformer Module 10(text)/10(audio) 0.1/50
Qwen 5,20,10,0.88 layer 26 of Transformer Module 10(text)/10(audio) 1.0/50
AF 5,20,10,0.00 layer 26 of Transformer Module 10(text)/10(audio) 1e-2/100
@@
FT (LLM) DeSTA - 5h 50m 2h 00m
Qwen - 2h 35m 1h 15m
AF - 2h 00m 0h 40m
FT (Audio) DeSTA - 4h 20m 2h 00m
Qwen - 3h 00m 1h 30m
AF - 1h 30m 0h 40m
KE DeSTA 4h 10m 8h 40m 9h 45m
Qwen 2h 10m 4h 50m 3h 20m
AF 2h 10m 3h 00m 2h 15m
MEND DeSTA 4h 10m 4h 00m 4h 10m
Qwen 2h 20m 2h 10m 1h 50m
AF 2h 20m 2h 00m 0h 25m
UnKE DeSTA - 3h 00m 1h 05m
Qwen - 1h 35m 0h 30m
AF - 1h 45m 0h 25m
I-IKE DeSTA - 6h 00m 3h 45m
Qwen - 3h 00m 2h 00m
AF - 1h 20m 0h 45m
IE-IKE DeSTA - 3h 20m 8h 00m
Qwen - 2h 00m 3h 50m
AF - 2h 10m 4h 00m
WISE DeSTA - 6h 10m 2h 00m
Qwen - 3h 00m 1h 10m
AF - 7h 00m 1h 45m
D.1 Editing Methods
Fine-tuning adapts the model via gradient descent on selected components. We apply it to the last layer of the LLM backbone (FT (LLM)) and the modality connector between audio encoders and the backbone in LALMs (FT (Audio)).
Knowledge Editor (KE) (De Cao et al., 2021) employs a hypernetwork to update parameters. It leverages a bidirectional LSTM with constrained optimization, in which generality and locality data is involved, to predict weight updates.
MEND (Mitchell et al., 2022) uses a hypernetwork to generate parameter updates by decomposing fine-tuning gradients into low-rank forms and transforming them into parameter updates. Similarly, it also leverages generality and locality data during training.
WISE (Wang et al., 2024b) performs knowledge editing by introducing a dual-memory architecture that separates the pretrained FFN weights from a lightweight side-memory used to store edit-specific updates. During editing, WISE optimizes a routing-aware activation loss that forces edited queries to rely on the side-memory while keeping irrelevant inputs mapped to the original parameters, ensuring strong locality without degrading pretrained behavior.
D.2 Implementation Details
The hyperparameters of each editing method are shown in Table 6, and their approximate execution time on NVIDIA H100 GPU is reported in Table 7. The execution time varies across models due to differences in their response patterns. DeSTA generally produces detailed and descriptive explanations, whereas AF typically generates concise answers without additional explanation. Qwen occasionally provides explanations, but this behavior is not consistent.
For FT (LLM) and FT (Audio), early stopping is applied when the loss falls below 1e-2.
For KE and MEND, which require training a hypernetwork, we train the hypernetwork for one epoch due to its rapid convergence on the validation set.
For UnKE, the editing process consists of two stages with separate hyperparameters. The first stage uses v_step iterations to find the desired key vector, with its own learning rate v_LR. The second stage uses optim_step iterations to update the weights of the chosen layer so that it naturally produces the new key vector, controlled by a separate learning rate optim_LR. To prevent interference with unrelated knowledge, stage 2 also incorporates preservation data. Specifically, we use 20 preservation samples for text and 5 for audio. Preservation data constrains the key generator to retain original key representations while inserting new ones, preventing global parameter drift.
For WISE, the edit copy is optimized with an edit target loss and an activation-distance margin loss, where $\alpha$ constrains out-of-scope drift, $\beta$ enforces sufficient in-scope change, and $\gamma$ separates in-scope and out-of-scope activations. At inference time, an activation-based gating rule using stored edit statistics and $\mathtt{act\_ratio}$ selects between original and edited weights, ensuring edits are triggered only for in-scope inputs via preservation data.
For I-IKE and IE-IKE, the edited knowledge is encoded via natural language instructions in the system prompt, following the prompt template provided below. In IE-IKE, demonstration examples are retrieved from the training set based on the editing pair. For example, when editing the model’s perception of speaker emotion from happy to sad, we retrieve an editing instance from the training set involving the same (happy, sad) editing pair. The demonstration examples illustrate how the model should respond to different types of evaluation queries, covering reliability, generality, and locality. Specifically, for each test sample, we provide 1 reliability example, 3 generality examples, 4 audio locality examples (or 3 when editing the speaker gender attribute), and 1 text locality example, all presented to the model as a multi-turn dialogue context.
System Prompt Template for I-IKE and IE-IKE You are performing in-context knowledge editing, but only for questions about audio inputs.
From now on, when analyzing any audio clip, you must apply the following edits consistently:
EDITS (New facts):
{edits}
where each entry has the form pre_edit $\rightarrow$ post_edit.
Rules:
1. If your reasoning or prediction about an audio clip would normally lead to pre_edit, you must instead treat it as post_edit.
2. All properties, attributes, and facts that belong to post_edit must be applied consistently, as if the audio were actually from post_edit.
3. If the user’s question is unrelated to these edits, you should answer normally without making changes.
4. Always ensure your final answers are fully consistent with the edited mapping.
In WISE, $\alpha$ bounds locality activation, with smaller values enforcing stricter preservation. $\beta$ sets a minimum edit activation, where larger values drive more aggressive edits. $\gamma$ enforces a margin ensuring edit activations exceed locality activations. Finally, act_ratio scales the inference routing threshold: higher values favor original weights, while lower values route more queries to the side-memory.
Appendix E More Details and Evaluation Prompts for LLM-as-a-judge
@@
We use GPT-5 mini (gpt-5-mini-2025-08-07) with minimal reasoning effort and a fixed random seed (0) as the judge in our LLM evaluator, balancing strong alignment with human judgments and cost efficiency. During evaluation, we focus on correctness and consistency. For reliability, generality, and portability, we assess whether the edited model’s responses correctly align with the ground truth. For locality, we assess whether the edited model remains consistent with the original model’s outputs. The evaluation prompts are provided below. To further reduce costs, some cases are judged directly without invoking the LLM evaluator: empty outputs are marked as incorrect/inconsistent, and responses that exactly match the ground truth (for correctness) or the original model’s output (for consistency) are marked as correct/consistent.
Human verification was conducted on 210 randomly selected samples for correctness and consistency, respectively, yielding 98.57% and 97.62% agreement, indicating the robustness of the employed LLM-as-a-judge.
@@
Ground truth answer: {ground_truth}
Model-generated response: {model_response}
Return format.
@@
Original Output: {original_output}
New Output: {new_output}
Task. Decide if these two outputs are semantically consistent. Outputs are consistent if they express the same main message, even with different wording, phrasing, or level of detail. Outputs are inconsistent if they contradict each other, reach different conclusions, or provide fundamentally different answers. Return format.
@@
Appendix F Detailed Results for Each Auditory Attribute under Single Editing
Table 8: Detailed results of the four metrics of each auditory attribute across different editing methods on DeSTA under single editing. Attr. denotes auditory attributes, and Port. denotes portability. For generality and audio locality, Avg. indicates the average performance across all types of the corresponding metric. (%)
@@ Appendix F Detailed Results for Each Auditory Attribute under Single Editing
FT (LLM) ALL 99.75 98.75 99.67 99.00 97.58 64.93 87.92 15.78 74.67 69.08 82.67 18.83
Animal 99.33 99.11 99.67 99.33 98.33 55.08 91.00 9.33 56.33 63.67 83.33 20.00
Emotion 100.00 99.56 99.33 100.00 99.33 53.17 80.67 7.33 59.67 65.00 77.67 22.00
Gender 99.67 96.78 99.67 96.67 94.00 84.44 90.33 - 88.33 74.67 88.33 6.00
Language 100.00 99.56 100.00 100.00 98.67 71.92 89.67 30.67 94.33 73.00 81.33 27.33
FT (Audio) ALL 99.50 86.06 96.75 84.83 76.58 68.13 78.58 48.11 76.00 64.83 100.00 55.33
Animal 99.00 85.67 97.33 83.00 76.67 62.08 81.67 40.67 60.00 66.00 100.00 29.33
Emotion 99.67 78.56 95.33 73.67 66.67 60.75 77.67 38.67 66.00 60.67 100.00 46.00
Gender 99.67 92.22 99.33 92.33 85.00 77.78 78.33 - 87.33 67.67 100.00 82.67
Language 99.67 87.78 95.00 90.33 78.00 74.33 76.67 65.00 90.67 65.00 100.00 63.33
KE ALL 99.58 99.17 99.25 99.25 99.00 79.49 96.42 43.89 76.33 92.42 93.08 17.33
Animal 98.67 97.33 97.67 97.33 97.00 77.42 97.00 61.67 59.00 92.00 93.00 23.33
Emotion 100.00 100.00 100.00 100.00 100.00 66.25 98.33 16.33 59.33 91.00 93.67 16.00
Gender 99.67 99.56 99.33 99.67 99.67 94.22 96.67 - 92.33 93.67 93.33 7.67
Language 100.00 99.78 100.00 100.00 99.33 83.75 93.67 53.67 94.67 93.00 92.33 22.33
MEND ALL 95.25 94.94 95.83 95.08 93.92 71.44 93.50 17.22 74.58 86.92 92.00 18.08
Animal 90.00 89.22 91.33 88.67 87.67 63.50 93.67 19.33 57.00 84.00 92.33 20.33
Emotion 99.67 98.44 99.00 99.00 97.33 62.17 94.67 9.00 59.33 85.67 93.33 23.67
Gender 98.67 97.00 97.33 99.33 94.33 91.78 97.00 - 88.33 90.00 90.33 4.67
Language 92.67 95.11 95.67 93.33 96.33 73.42 88.67 23.33 93.67 88.00 92.00 23.67
UnKE ALL 96.09 87.42 87.92 92.50 81.84 57.18 67.59 10.78 72.25 62.83 88.50 17.17
Animal 99.67 96.11 98.00 99.67 90.67 52.75 69.67 16.33 60.00 65.00 88.33 27.33
Emotion 95.00 87.67 87.33 90.67 85.00 50.50 66.67 12.00 57.33 66.00 87.67 18.33
Gender 99.67 86.11 86.00 99.67 72.67 71.89 58.33 - 94.33 63.00 88.00 6.00
Language 90.00 79.78 80.33 80.00 79.00 53.58 75.67 4.00 77.33 57.33 90.00 17.00
I-IKE ALL 64.25 61.08 64.67 58.83 59.75 65.49 79.33 50.67 73.67 54.58 61.92 71.33
Animal 90.33 92.67 92.67 92.67 92.67 54.92 83.67 22.00 59.67 54.33 60.67 78.33
Emotion 59.00 49.78 58.00 44.67 46.67 59.92 80.67 45.33 62.00 51.67 59.67 65.33
Gender 69.00 64.11 71.00 60.33 61.00 73.33 81.33 - 79.67 59.00 63.67 72.67
Language 38.67 37.78 37.00 37.67 38.67 75.75 71.67 84.67 93.33 53.33 63.67 69.00
IE-IKE ALL 40.58 39.61 41.50 38.67 38.67 58.89 70.00 50.00 70.75 42.58 56.67 34.33
Animal 77.00 78.56 80.00 76.67 79.00 51.33 78.67 23.33 59.00 44.33 60.67 44.00
Emotion 27.00 26.67 28.67 25.33 26.00 51.67 68.00 33.67 61.33 43.67 52.33 28.67
Gender 41.67 37.67 43.00 36.00 34.00 61.56 74.33 - 67.67 42.67 60.67 32.67
Language 16.67 15.56 14.33 16.67 15.67 71.67 59.00 93.00 95.00 39.67 53.00 32.00
WISE ALL 100.00 100.00 100.00 100.00 100.00 38.90 29.52 3.78 76.00 31.38 83.08 17.34
Animal 100.00 100.00 100.00 100.00 100.00 33.94 39.33 5.00 58.00 33.44 83.33 11.33
Emotion 100.00 100.00 100.00 100.00 100.00 24.17 16.33 6.33 56.33 17.67 76.00 2.68
Gender 100.00 100.00 100.00 100.00 100.00 63.44 43.67 - 93.67 53.00 90.00 29.67
Language 100.00 100.00 100.00 100.00 100.00 34.06 18.73 0.00 96.00 21.40 83.00 25.67
@@
Table 9: Detailed results of the four metrics of each auditory attribute across different editing methods on Qwen under single editing. Attr. denotes auditory attributes, and Port. denotes portability. For generality and audio locality, Avg. indicates the average performance across all types of the corresponding metric. (%)